OpenStudy (anonymous):

Find a polynomial f(x) with integer coefficients such that one root f(x) = 0 is x = √2 + ³√2

OpenStudy (anonymous):

I don't know how to go about solving this problem, can anyone show me the steps?

OpenStudy (bahrom7893):

Okay, uhmmm x1 = (-b - sqrt(D))/(2a), x2 = (-b + sqrt(D))/(2a)

OpenStudy (bahrom7893):

Where D = b^2 - 4ac so you have: x2 = (-b + sqrt(b^2-4ac))/(2a)

OpenStudy (bahrom7893):

Or, x2 = (-b/(2a)) + sqrt(b^2 - 4ac)/(2a)

OpenStudy (bahrom7893):

Which means that (-b/(2a)) = sqrt(2) and sqrt(b^2-4ac)/(2a) = cuberoot(2)

OpenStudy (bahrom7893):

-b = 2a*sqrt(2) sqrt(8a^2 - 4ac)/(2a) = cuberoot(2) (8a^2-4ac)/(4a^2) = 2^(2/3) (2a - c) = 2^(2/3) a

OpenStudy (bahrom7893):

We need to like plug in something for c or a, i think. i gotta go now, @satellite73 @amistre64 @experimentX @estudier can one of you finish this? Thanks :)

OpenStudy (anonymous):

Based on your work, you're assuming its a quadratic. Why are you allowed to make that assumption? Since there is a cube root, wouldn't it be a third degree polynomial?

OpenStudy (experimentx):